### v1

# Distributed Omniscient Observers for Multi-Agent Systems

Abstract
This article proposes fully distributed omniscient observers for both heterogeneous and homogeneous linear multi-agent systems, through which each agent can estimate the states of all agents. The proposed observers not only contribute to distributed Nash equilibrium seeking in multi-player games, but also provide a designable self-organization mechanism for artificial swarms to emulate biological social behaviors, including sheepdog herding and honeybee dance communication.
I Introduction
Consensus is widely recognized as one of the most fundamental cooperative behaviors in multi-agent systems (MAS). Analogous to those biological synchronies (e.g., synchronous flashing of fireflies [2]), it typically describes a phenomenon that the state trajectories of all agents evolve identically.
Since the pioneering research in [5, 14, 25], consensus control of multi-agent systems has been extensively investigated over the past two decades. Most research formulates the consensus problem within a distributed context, where each agent only has access to limited local information and can only communicate with neighboring agents [24, 18]. Numerous consensus protocols have been developed under various complications, including model uncertainties [17, 7], switching and disturbed communication links [23, 16, 28], velocity and acceleration constraints [19, 8, 9, 6], etc.
As consensus-reaching has been relatively well studied, recent focus has been increasingly placed on more advanced forms of cooperation in MAS. However, the aforementioned distributed setup poses challenges to achieving advanced collaboration. Specifically, if an agent only has access to limited local information, it may fail to effectively cooperate with others for a global objective of the MAS. This motivates the development of distributed omniscient observers in this article, which aim to provide each agent with sufficient global information to support autonomous decision-making.
In prior research on leader–follower MAS, distributed observers have been commonly designed either for each follower to estimate the leader’s state [11, 13, 20, 26, 27, 21, 22], or for each agent to reconstruct its own absolute state using relative output information [29]. Different from the above approaches, the distributed omniscient observers proposed in this article enable every agent to estimate the global MAS state (which includes both its own state and the states of all other agents), while allowing each agent’s input to be persistently nonzero. This equips each agent with a “global view”, enabling the MAS to cooperatively perform complex tasks that extend beyond basic consensus. The proposed observer design method mainly builds on the framework developed in [4]. However, the method in this article emphasizes exploiting relative (neighbor-to-neighbor) output information in MAS, which can be easier and cheaper to realize in practice compared with using the absolute one.
Notation
[…]
II Review of the Distributed Observers Without Using Global Inputs
II-A Communication Graph
The communication links among observer nodes in distributed observers (or among agents in MAS) enable information exchange. The topology of the communication links can be characterized using a bidirected graph, introduced as follows.
A graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ consists of a finite, nonempty node set $\mathcal{N}=\left\{1,2,\cdots,N\right\}$, and an edge set $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$, whose elements are ordered pairs of nodes. An edge originating from node $j$ and ending at node $i$ is denoted by $(j,i)\in\mathcal{E}$, which represents a directed information flow from node $j$ to node $i$. The adjacency matrix of $\mathcal{G}$ is denoted by $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$, where $a_{ij}$ denotes the weight of the corresponding edge, with $a_{ij}>0$ if $(j,i)\in\mathcal{E}$, and $a_{ij}=0$ otherwise. We assume that the graph has no self-loops, i.e., $a_{ii}=0$, $\forall i\in\mathcal{N}$. The Laplacian matrix $\mathcal{L}=[l_{ij}]\in\mathbb{R}^{N\times N}$ of graph $\mathcal{G}$ is defined by $l_{ii}=\sum\nolimits_{k=1}^{N}{{a_{ik}}}$ and $l_{ij}=-a_{ij},\ \forall i,j\in\mathcal{N},\ i\neq j$. A directed path from node $i$ to node $j$ is a sequence of edges $(i_{k-1},\ i_{k})\in\mathcal{E},\ k=1,2,\cdots,\bar{k}$, where $i_{0}=i,\ i_{\bar{k}}=j$. Graph $\mathcal{G}$ is said to be bidirected if $a_{ij}=a_{ji},\forall i,j\in\mathcal{N}$. A bidirected graph is said to be connected if, for every pair of distinct nodes $i,j\in\mathcal{N},\ i\neq j$, there exists at least one directed path from node $i$ to node $j$.
II-B Revisiting the Design Method
In [4], a distributed observer design method was proposed for the following linear time-invariant system:
$\dot{x}$ $=Ax+Bu$ (1a)
$y_{i}$ $=C_{i}x,\ i\in\mathcal{N},$ (1b)
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{m}$, and $y_{i}\in\mathbb{R}^{p_{i}}$ are the system state, input, and local output, respectively; $C_{i}$ has full row rank. In brief, the first step of the design procedure is to find matrices ${T_{id}}\in\mathbb{R}^{n\times\delta_{i}}$, ${B_{i}}\in\mathbb{R}^{n\times m_{i}}$, and ${B_{-i}}\in\mathbb{R}^{n\times m_{-i}}$ such that
$\sum\nolimits_{i=1}^{N}{{\mathop{\rm Im}\nolimits}{T_{id}}}$ $=\mathbb{R}^{n}$ (2a)
$\operatorname{Im}B_{i}+\operatorname{Im}B_{-i}$ $=\operatorname{Im}B$ (2b)
$T_{id}^{\top}{T_{id}}$ $=I_{\delta_{i}}$ (2c)
$m_{i},\ m_{-i}\leq m{\rm{\ and\ }}\delta_{i}$ $\leq n,\ \forall i\in\mathcal{N},$ (2d)
and meanwhile, find matrices ${E_{i}}\in{\mathbb{R}^{\delta_{i}\times\delta_{i}}}$, ${F_{i}}\in{\mathbb{R}^{\delta_{i}\times{p_{i}}}}$, and $G_{i}\in{\mathbb{R}^{\delta_{i}\times{p_{i}}}}$ that satisfy either
${G_{i}}{C_{i}}{B_{-i}}$ $=T_{id}^{\top}{B_{-i}}$ (2e)
${E_{i}}T_{id}^{\top}+({F_{i}}-{E_{i}}{G_{i}}){C_{i}}$ $=(T_{id}^{\top}-{G_{i}}{C_{i}})A$ (2f)
${\rm{Re}}\lambda({E_{i}})$ $<0,$ (2g)
or
$\delta_{i}$ $=p_{i}$ (2h)
$E_{i}=F_{i}$ $=0_{p_{i}\times p_{i}}$ (2i)
$G_{i}C_{i}$ $=T_{id}^{\top},$ (2j)
for each $i\in\mathcal{N}$. According to (2b), there exist $u_{i}(t)$ and $u_{-i}(t)$ such that [^1] In [4], $u_{i}$ and $u_{-i}$ denote available and unavailable input information to the $i$th observer node, respectively.
$Bu(t)=B_{i}u_{i}(t)+B_{-i}u_{-i}(t),\ \forall i\in\mathcal{N},\ \forall t\geq 0.$ (5)
Based on (2c), there exists $T_{iu}$ such that
$T_{id}^{\top}T_{iu}=0\ {\rm{and}}\ T_{iu}^{\top}T_{iu}=I_{n-\delta_{i}},\ \forall i\in\mathcal{N}.$
The dynamics of the $i$th observer node are of the following form:
${{\dot{z}}_{i}}$ $={{\bar{E}}_{i}}{z_{i}}+{{\bar{F}}_{i}}{y_{i}}+{{\bar{B}}_{i}}{u_{i}}-{H_{i}}\left[{\sum\nolimits_{j=1}^{N}{{a_{ij}}({{\hat{x}}_{i}}-{{\hat{x}}_{j}})}}\right]$ (6a)
${{\hat{x}}_{i}}$ $={z_{i}}+{{\bar{G}}_{i}}{y_{i}},$ (6b)
where ${z}_{i}$, with initial value ${z}_{i}(0)=0$, is an intermediate variable; ${\hat{x}}_{i}$ is the estimate of state $x$, produced by the $i$th observer node; $u_{i}$ is from (5). The matrix gains in (6) are designed as
${{\bar{E}}_{i}}$ $={T_{id}}{E_{i}}T_{id}^{\top}+{T_{iu}}T_{iu}^{\top}A$ (7a)
${{\bar{F}}_{i}}$ $={T_{id}}{F_{i}}+{T_{iu}}T_{iu}^{\top}A{{\bar{G}}_{i}}$ (7b)
${{\bar{G}}_{i}}$ $={T_{id}}{G_{i}}$ (7c)
${{\bar{B}}_{i}}$ $=(I-{\bar{G}}_{i}{C_{i}}){B_{i}}.$ (7d)
For notational simplicity, let $\varepsilon_{iu}=T_{iu}^{\top}\sum\nolimits_{j=1}^{N}{a_{ij}(\hat{x}_{i}-\hat{x}_{j})}$, and define a vector function ${h}(\cdot)$ as
${h}(\omega)=\left\{\begin{aligned} \left\|\omega\right\|^{-1}\omega,\ \omega&\neq 0\\ 0,\ \omega&=0.\end{aligned}\right$.
The function $H_{i}(\cdot)$ in (6) is designed as follows:
${H_{i}}(\cdot)={\gamma_{i}}{T_{iu}}\varepsilon_{iu}+{\gamma_{is}}{T_{iu}}{h}(\varepsilon_{iu}),$ (8)
where $\gamma_{i}$ and $\gamma_{is}$ are scalar gains evolving according to the following adaptive laws:
${{\dot{\gamma}}_{i}}$ $={\phi_{i}}{\left\|\varepsilon_{iu}\right\|^{2}}$
${{\dot{\gamma}}_{is}}$ $={\phi_{is}}\left\|\varepsilon_{iu}\right\|$
with step sizes $\phi_{i}$}, $\phi_{is}$} and initial values $\gamma_{i}(0)$, $\gamma_{is}(0)$ chosen as positive constants. The convergence of the state estimation errors is guaranteed under the following assumptions:
Assumption 1.
There is a finite bound for $u_{-i}(t)$ in (5), i.e., $\exists\bar{u}_{-}\in\mathbb{R},\ {\rm{s.t.}}\ \mathop{\max}\nolimits_{t\geq 0}\left\|u_{-i}(t)\right\|\leq\bar{u}_{-},\ \forall i\in\mathcal{N}$.
Assumption 2.
Communication graph $\mathcal{G}$ is connected.
One of the main results in [4] is as follows:
Lemma 1.
[4] Under Assumptions 1 and 2, the distributed observers with node dynamics (6) can produce accurate state estimates for system (1), i.e.,
$\mathop{\lim}\nolimits_{t\to\infty}\left\|{{{\hat{x}}_{i}}(t)-x(t)}\right\|=0,\ \forall i\in\mathcal{N}.$
Moreover, adaptive gains $\gamma_{i}$ and $\gamma_{is}$ remain bounded, $\forall i\in\mathcal{N}$.
III Distributed Omniscient Observer Design
Given matrices $A$, $B_{-i}$, and $C_{i}$, an algorithm in [4] can compute a numerical solution $\left\{T_{id},E_{i},F_{i},G_{i}\right\}$ that satisfies (2c)-(2g), or satisfies (2c), (2d), and (2h)-(2j). In this section, we show that a block-diagonal structure of matrices $A$ and $B$ in (1a) admits analytical solutions $\left\{B_{i},B_{-i},T_{id},E_{i},F_{i},G_{i}\mid i\in\mathcal{N}\right\}$ that satisfy the constraints posed in (2). These analytical solutions are useful in the design of distributed omniscient observers for MAS.
III-A Design for Heterogeneous Multi-Agent Systems
Consider a group of $N$ agents that have heterogeneous, general linear dynamics. The dynamics of the $i$th agent are described by
$\dot{\breve{x}}_{i}$ $=\breve{A}_{i}\breve{x}_{i}+\breve{B}_{i}u_{i}$ (9a)
${y}_{i}$ $=\breve{C}_{i}\breve{x}_{i},\ i\in\mathcal{N},$ (9b)
[…]
If $(\breve{A}_{i},\breve{C}_{i})$ is detectable for all $i\in\mathcal{N}$, then an analytical solution that satisfies (2) is as follows:
[…]
$E_{i}$ $=\breve{A}_{i}+\breve{L}_{i}\breve{C}_{i},\ {F_{i}}=-{\breve{L}_{i}},\ G_{i}=0_{n_{i}\times p_{i}},\ i\in\mathcal{N},$ (10i)
where $\breve{L}_{i}$ is chosen such that $\breve{A}_{i}+\breve{L}_{i}\breve{C}_{i}$ has eigenvalues with negative real parts. With $u_{-i}$ chosen as the inputs of all agents except agent $i$, the fact that $u_{i}$ is the input of the $i$th agent can preserve (5). Then the following result is straightforward from Lemma 1.
Theorem 1.
Consider heterogeneous MAS (9), where $(\breve{A}_{i},\breve{C}_{i})$ is detectable and $u_{i}$ is bounded, $\forall i\in\mathcal{N}$. Suppose the agents can collect their own input $u_{i}$ and output $y_{i}$, and implement observer dynamics (6) over a connected communication graph $\mathcal{G}$. If the observers are designed according to (7), (8), and (10), then each agent can produce an accurate global state estimate, i.e.,
$\mathop{\lim}\nolimits_{t\to\infty}\left\|{{{\hat{x}}_{i}}(t)-x(t)}\right\|=0,\ \forall i\in\mathcal{N}.$
Moreover, adaptive gains $\gamma_{i}$ and $\gamma_{is}$ remain bounded, $\forall i\in\mathcal{N}$.
The distributed omniscient observers given in Theorem 1 enable each agent to estimate the global state of MAS (9), by only using its own input and output information while exchanging estimates with neighboring agents.
III-B Design for Homogeneous Multi-Agent Systems
It follows directly that the distributed omniscient observers developed in Section III-A can be applied to the following homogeneous MAS:
$\dot{\breve{x}}_{i}$ $=\breve{A}\breve{x}_{i}+\breve{B}\breve{u}_{i}$ (11a)
$\breve{y}_{i}$ $=\breve{C}\breve{x}_{i},\ i\in\mathcal{N},$ (11b)
where $\breve{x}_{i}\in\mathbb{R}^{\breve{n}}$ is the state, $\breve{u}_{i}\in\mathbb{R}^{\breve{m}}$ is the control input, and $\breve{y}_{i}\in\mathbb{R}^{\breve{p}}$ is the measured output. However, this requires each agent to measure its own output. For homogeneous MAS, the remainder of this section addresses the case where most agents can only measure the relative output between themselves and neighboring agents.
Let $\mathcal{R}$ denote a proper subset of $\mathcal{N}$, formed by the indices of those agents that have access to their own outputs. Accordingly, define the following two variables:
${u}_{i}=$ $\begin{cases}\sum\nolimits_{j=1}^{N}{{a_{ij}}(\breve{u}_{i}-\breve{u}_{j})},&i\in\mathcal{N}\setminus\mathcal{R}\\ \ \ \ \ \ \ \ \ \ \breve{u}_{i},&i\in\mathcal{R}\end{cases}$ (12a)
${y}_{i}=$ $\begin{cases}\sum\nolimits_{j=1}^{N}{{a_{ij}}(\breve{y}_{i}-\breve{y}_{j})},&i\in\mathcal{N}\setminus\mathcal{R}\\ \ \ \ \ \ \ \ \ \ \breve{y}_{i},&i\in\mathcal{R},\end{cases}$ (12b)
which should be collected by the $i$th agent and be fed into observer dynamics (6). Let $\varpi_{i}$} denote the $i$th column vector of identity matrix $I_{N}$, and define the following row vector:
$\bar{\mathcal{L}}_{i}=\begin{cases}\varpi_{i}^{\top}\mathcal{L},&i\in\mathcal{N}\setminus\mathcal{R}\\ \ \varpi_{i}^{\top},&i\in\mathcal{R}.\end{cases}$ (13)
The dynamics of the overall MAS take the same form as (1a), where $x={\rm{col}}(\breve{x}_{i})_{i=1}^{N}$, $u={\rm{col}}(\breve{u}_{i})_{i=1}^{N}$, $A=I_{N}\otimes\breve{A}$, and $B=I_{N}\otimes\breve{B}$. Moreover, (12b) aligns with (1b), where $C_{i}=\breve{C}\left({{\bar{\cal L}}_{i}}\otimes{I_{\breve{n}}}\right)$. In this case, we construct an analytical solution of the following form:
${B_{i}}$ $=\frac{{T_{id}}{\breve{B}}}{\left\|{{{\bar{\cal L}}_{i}}}\right\|}$ (14a)
${B_{-i}}$ $=\left({{I_{\breve{n}N}}-{T_{id}}T_{id}^{\top}}\right)\left({I_{N}}\otimes\breve{B}\right)$ (14b)
$T_{id}^{\top}$ $=\frac{{{\bar{\cal L}}_{i}}\otimes I_{\breve{n}}}{\left\|{{\bar{\cal L}}_{i}}\right\|}$ (14c)
$E_{i}$ $=\breve{A}+\breve{L}_{i}\breve{C},\ G_{i}=0_{\breve{n}\times\breve{p}}$ (14d)
${F_{i}}$ $=-\frac{{\breve{L}_{i}}}{\left\|{{{\bar{\cal L}}_{i}}}\right\|},\ i\in\mathcal{N},$ (14e)
where ${\bar{\cal L}}_{i}$ is defined in (13), and $\breve{L}_{i}$ is chosen such that $\breve{A}+\breve{L}_{i}\breve{C}$ has eigenvalues with negative real parts. It can be verified that the above analytical solution satisfies constraints (2c)-(2g). The following two lemmas guarantee that constraints (2a) and (2b) are also satisfied.
Lemma 2.
If $\mathcal{L}\in\mathbb{R}^{N\times N}$ is a Laplacian matrix of a connected graph and $\mathcal{R}\neq\emptyset$, then $\bar{\mathcal{L}}_{i}$ defined in (13) forms a nonsingular matrix ${\rm{col}}(\bar{\mathcal{L}}_{i})_{i=1}^{N}$.
Lemma 3.
With $T_{id}$ selected as in (14c), the column spaces of matrices ${T_{id}}\breve{B}$ and ${T_{id}}T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)$} are identical, i.e., ${\rm{Im}}\left({{T_{id}}\breve{B}}\right)={\rm{Im}}\left[{{T_{id}}T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)}\right]$.
See Sections VII-A and VII-B for the proof of Lemma 2 and Lemma 3, respectively. Given the fact that choosing $u_{i}$ as in (12a) and $u_{-i}$ as $u$ preserves (5), the following result is straightforward from Lemma 1.
Theorem 2.
Consider homogeneous MAS (11), where $(\breve{A},\breve{C})$ is detectable and $\breve{u}_{i}$ is bounded, $\forall i\in\mathcal{N}$. Suppose the agents can collect input-output information (12) and implement observer dynamics (6) over a connected communication graph $\mathcal{G}$. If $\mathcal{R}\neq\emptyset$ and the observers are designed according to (7), (8), and (14), then each agent can produce an accurate global state estimate, i.e.,
$\mathop{\lim}\nolimits_{t\to\infty}\left\|{{{\hat{x}}_{i}}(t)-x(t)}\right\|=0,\ \forall i\in\mathcal{N}.$ (15)
Moreover, adaptive gains $\gamma_{i}$ and $\gamma_{is}$ remain bounded, $\forall i\in\mathcal{N}$.
Provided that there is at least one agent having access to its own output, the distributed omniscient observers given in Theorem 2 enable each agent to estimate the global state of MAS (11) by using relative input-output information. Moreover, the observer gain design of the $i$th agent only relies on the dynamic model of the agents and the $i$th row vector of the Laplacian matrix; it does not require knowledge of the full Laplacian matrix.
III-C Further Extension
To implement the observers designed in Section III-B, each agent needs neighbors’ input information. In what follows, we develop distributed omniscient observers that do not require neighbors’ inputs for homogeneous MAS (11). This is achieved at the expense of undermining adaptability to the change of communication graphs, since a unified scalar gain depending on the global graph should be shared by each agent. Specifically, let us consider the distributed observers proposed in [29]:
$\dot{\hat{\breve{x}}}_{i}=\breve{A}{\hat{\breve{x}}}_{i}+\breve{B}\breve{u}_{i}+cM\zeta_{i},\ i\in\mathcal{N},$ (16)
where ${\hat{\breve{x}}}_{i}$ is the estimate of ${\breve{x}}_{i}$, and $\zeta_{i}$ is designed as
$\zeta_{i}=w_{i}\left({\breve{y}}_{i}-\breve{C}{\hat{\breve{x}}}_{i}\right)+{\sum\nolimits_{j=1}^{N}{{a_{ij}}\left[{\breve{y}}_{i}-{\breve{y}}_{j}-\breve{C}\left({\hat{\breve{x}}}_{i}-{\hat{\breve{x}}}_{j}\right)\right]}}$ (17)
with $w_{i}>0$ if the $i$th agent has access to its own output ${\breve{y}}_{i}$, and $w_{i}=0$ otherwise. The scalar gain $c$ and matrix gain $M$ in (16) are designed as follows:
$c\geq\frac{1}{2\lambda_{\min}\left(\mathcal{L}+W\right)},\ M=S\breve{C}^{\top},$ (18)
where $W={\rm{diag}}(w_{i})_{i=1}^{N}$, and $S$ is the unique positive definite solution of
$\breve{A}S+S\breve{A}^{\top}-S\breve{C}^{\top}\breve{C}S+I=0.$
Based on (16), the design of distributed omniscient observers (6) is given as follows:
$A$ $=I_{N}\otimes\breve{A},\ B_{i}=0_{\breve{n}N\times 1},\ u_{i}=0,\ y_{i}={\hat{\breve{x}}}_{i}$ (19a)
$T_{id}^{\top}$ $=C_{i}=\begin{bmatrix}0_{\breve{n}\times(i-1)\breve{n}}&I_{\breve{n}}&0_{\breve{n}\times(N-i)\breve{n}}\end{bmatrix}$ (19b)
$E_{i}$ $=F_{i}=0_{\breve{n}\times\breve{n}},\ G_{i}=I_{\breve{n}},\ \forall i\in\mathcal{N}.$ (19c)
Theorem 3.
Consider homogeneous MAS (11), where $(\breve{A},\breve{C})$ is detectable and $\breve{u}_{i}$ is bounded, $\forall i\in\mathcal{N}$. Suppose the agents implement observer dynamics (6) and (16) over a connected communication graph $\mathcal{G}$. If $W\neq 0$ and the observers are designed according to (7), (8), and (17)-(19), then each agent can produce an accurate global state estimate. Moreover, adaptive gains $\gamma_{i}$ and $\gamma_{is}$ remain bounded, $\forall i\in\mathcal{N}$.
See Section VII-C for the proof of Theorem 3.
IV Application A: Distributed Nash Equilibrium Seeking in Multi-Player Games
Consider a set of players indexed from $1$ to $N$. For each player $i\in\mathcal{N}$, let $\breve{x}_{i}\in\mathbb{R}^{\breve{n}}$ denote its action, and $J_{i}(x):\mathbb{R}^{\breve{n}N}\rightarrow\mathbb{R}$ denote its cost function, where $x={\rm{col}}(\breve{x}_{i})_{i=1}^{N}$. Define $\breve{x}_{-i}=\begin{bmatrix}\breve{x}_{1}&\cdots&\breve{x}_{i-1}&\breve{x}_{i+1}&\cdots&\breve{x}_{N}\end{bmatrix}^{\top}$. The Nash equilibrium problem can be described as follows [12, 1]:
$\min\nolimits_{\breve{x}_{i}\in\mathbb{R}^{\breve{n}}}J_{i}\left(\breve{x}_{i},\breve{x}_{-i}\right),\ \forall i\in\mathcal{N}.$
Accordingly, the Nash equilibrium refers to an action profile of all players $x^{*}={\rm{col}}\left(x_{i}^{*}\right)_{i=1}^{N}$ that satisfies
$J_{i}\left(x_{i}^{*},x_{-i}^{*}\right)\leq J_{i}\left(\breve{x}_{i},x_{-i}^{*}\right),\ \forall\breve{x}_{i}\in\mathbb{R}^{\breve{n}},\ \forall i\in\mathcal{N}.$
At the Nash equilibrium, no player can diminish its own cost by unilaterally changing its action.
To solve the Nash equilibrium problem, define the game mapping as $\nabla J(x)={\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)\right]_{i=1}^{N}$, provided that the cost function $J_{i}$ is continuously differentiable in $\breve{x}_{i}$. The game mapping is said to be strongly monotone with constant $\mu>0$, if it holds that $\left(x_{a}-x_{b}\right)^{\top}\left[\nabla J(x_{a})-\nabla J(x_{b})\right]\geq\mu\left\|x_{a}-x_{b}\right\|^{2}$ for any $x_{a},x_{b}\in\mathbb{R}^{\breve{n}N}$. The following is a basic centralized Nash equilibrium seeking algorithm:
$\dot{\breve{x}}_{i}=-\nabla_{\breve{x}_{i}}J_{i}(x),\ \forall i\in\mathcal{N}.$ (20)
Lemma 4.
[12] Suppose that each cost function $J_{i}\left(\breve{x}_{i},\breve{x}_{-i}\right)$ is continuously differentiable and convex in $\breve{x}_{i}$ for every fixed $\breve{x}_{-i}$. If game mapping $\nabla J(x)$ is strongly monotone, then there exists a unique Nash equilibrium for the game and the trajectory of (20) converges to it, i.e.,
$\mathop{\lim}\nolimits_{t\to\infty}\left\|{{\breve{x}_{i}}(t)-x_{i}^{*}}\right\|=0,\ \forall i\in\mathcal{N}.$
Algorithm (20) requires each player to have real-time access to global action profile $x$. For the sake of scalability, however, communications may only occur between neighboring agents in MAS. In the case where only neighbors’ actions are directly available, the distributed omniscient observers developed in Section III can be used to implement algirithm (20), by providing each player with an estimate of the global action profile. Specifically, the distributed Nash equilibrium seeking algorithm is designed as follows:
$\dot{\breve{x}}_{i}={\breve{u}}_{i},\ \forall i\in\mathcal{N},$ (21)
where ${\breve{u}}_{i}=-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})$, and $\hat{x}_{i}$ is produced by the distributed omniscient observers presented in Theorem 1 or Theorem 2.
Theorem 4.
Suppose that each cost function $J_{i}\left(\breve{x}_{i},\breve{x}_{-i}\right)$ is continuously differentiable and convex in $\breve{x}_{i}$ for every fixed $\breve{x}_{-i}$. Moreover, suppose that there exist two constants $\chi,\chi_{s}\geq 0$ such that
$\left\|\nabla_{\breve{x}_{i}}J_{i}(x_{a})-\nabla_{\breve{x}_{i}}J_{i}(x_{b})\right\|^{2}\leq\chi\left\|x_{a}-x_{b}\right\|^{2}+\chi_{s}\left\|x_{a}-x_{b}\right\|,$ (22)
$\forall x_{a},x_{b}\in\mathbb{R}^{\breve{n}N},\ \forall i\in\mathcal{N}$. If the game mapping $\nabla J(x)$ is strongly monotone, then implementing algorithm (21) based on the observers given in Theorem 1 or Theorem 2 gives the unique Nash equilibrium of the game, i.e.,
$\mathop{\lim}\nolimits_{t\to\infty}\left\|{{\breve{x}_{i}}(t)-x_{i}^{*}}\right\|=0,\ \forall i\in\mathcal{N}.$
See Section VII-D for the proof of Theorem 4. Further discussions on Theorem 4 are as follows:
  • Condition (22) is more general than the following Lipschitz condition: There exists a constant $\bar{\chi}\geq 0$ such that
    $\left\|\nabla_{\breve{x}_{i}}J_{i}(x_{a})-\nabla_{\breve{x}_{i}}J_{i}(x_{b})\right\|\leq\bar{\chi}\left\|x_{a}-x_{b}\right\|,$ (23)
    $\forall x_{a},x_{b}\in\mathbb{R}^{\breve{n}N},\ \forall i\in\mathcal{N}$. For example, for scalar $\breve{x}_{i}$,
    $\nabla_{\breve{x}_{i}}J_{i}(x)=\begin{cases}\breve{x}_{i}+\sqrt{\breve{x}_{i}},&\breve{x}_{i}\geq 0\\ \breve{x}_{i}-\sqrt{-\breve{x}_{i}},&\breve{x}_{i}<0\end{cases}$
    satisfies (22), while it does not satisfy (23).
  • According to Theorem 1 and Theorem 2, if ${\breve{u}}_{i}$ in (21) is bounded [^2] Taking ${\breve{u}}_{i}=-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})$ as an example, the boundedness can be guaranteed by assuming that each player’s action belongs to a bounded closed subset of $\mathbb{R}^{\breve{n}}$, and $\nabla_{\breve{x}_{i}}J_{i}$ satisfies Lipschitz condition on this subset with a Lipschitz extension [10] outside this subset., the distributed omniscient observers can fulfill (15), which does not rely on the specific value of ${\breve{u}}_{i}$. This implies that the seeking algorithm and the distributed omniscient observers can be designed separately. The separability may help accommodate a variety of seeking algorithms for the solution of more complex Nash equilibrium problems in future research.
V Application B: Self-Organized Social Behavior Emergence in Artificial Swarms
Two bio-inspired simulation examples in this section demonstrate possible use cases of the proposed distributed omniscient observers. Since there is no command center coordinating the agents, the following decision and action mechanism is referred to as a self-organized way to bring out collective intelligent behaviors of them.
V-A Confine Companions to a Convex Hull
The first example is inspired by the herding behaviors of sheepdogs — they collaborate with each other to gather and move livestock from one place to another. In this example, there are leader agents and follower agents. The leaders can move freely, which represents the behavior of herding sheepdogs. The followers will identify which agents are leaders and assemble into the convex hull formed by the leaders.
Basic Setup: Within an x-y plane, dynamics of the agents indexed from $1$ to $N$ are of the form (11), where
${\breve{x}}_{i}=$ $\begin{bmatrix}{\breve{p}}_{i}^{x}\\ {\breve{p}}_{i}^{y}\\ {\breve{z}}_{i}\end{bmatrix},\ {\breve{A}}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},\ {\breve{B}}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$,
${\breve{u}}_{i}=$ $\begin{bmatrix}{\breve{v}}_{i}^{x}\\ {\breve{v}}_{i}^{y}\\ {\breve{u}}_{i}^{z}\end{bmatrix},\ {\breve{C}}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$,
with ${\breve{p}}_{i}^{x}$ and ${\breve{p}}_{i}^{y}$ denoting the positions, ${\breve{v}}_{i}^{x}$ and ${\breve{v}}_{i}^{y}$ the velocities, ${\breve{z}}_{i}$ the identity state, and ${\breve{u}}_{i}^{z}$ the identity input of the $i$th agent. Based on a connected graph $\mathcal{G}$ and a nonempty set $\mathcal{R}$, the agents collect relative/absolute input-output information, i.e., $u_{i}$ and $y_{i}$ defined in (12), and carry out the distributed omniscient observers presented in Theorem 2.
Leaders’ Actions: The velocities of leaders are freely chosen [^3] The trajectories of leaders can be designed by choosing their velocities, which can be used to guide followers through obstacles, or to serve other practical purposes.. The identity input of a leader is chosen as ${\breve{u}}_{j}^{z}=-{\breve{z}}_{j}+z^{*}$, where $j$ is the leader agent’s index and $z^{*}$ is a positive constant. The identity input is used to increase the identity state, so that followers can tell which agents are leaders according to the estimated identity states produced by the distributed omniscient observers.
Followers’ Decisions: Each follower determines in real time a set of candidate leaders for itself, based on the estimates of the identity states of all the other agents. From the perspective of a follower agent, anyone of the other agents will be labeled as a candidate leader, if the estimated identity state of the agent is greater than $z^{*}_{t}$, a positive threshold chosen to be lower than $z^{*}$.
Followers’ Actions: Each follower heads toward a candidate target point, that is a convex combination of the estimated positions of the candidate leaders. A follower with index $k$ will use the estimate for ${\breve{x}}_{k}$ provided in $\hat{x}_{k}$ to design control inputs ${\breve{v}}_{i}^{x}$ and ${\breve{v}}_{i}^{y}$ for itself. The identity input of the follower is chosen as ${\breve{u}}_{k}^{z}=-{\breve{z}}_{k}$.
Figure 1: Agent trajectories (start/end positions denoted by circles/crosses) in Section V-A.
V-B Summon Companions by Circling
The second example is inspired by the dancing behaviors of honeybees — they use dance language to communicate the location and the abundance of nectar sources to other members of the hive. In this example, the leader agents will circle at different places with different speeds, which simulates the behavior of the dancing honeybees. The follower agents will assemble at the leaders’ places respectively, according to the ratio of their circling speeds. Followers tend to be attracted to leaders with higher speeds.
Basic Setup: Within an x-y plane, dynamics of the agents indexed from $1$ to $N$ are of the form (11), where
${\breve{x}}_{i}=$ $\begin{bmatrix}{\breve{p}}_{i}^{x}\\ {\breve{p}}_{i}^{y}\\ {\breve{v}}_{i}^{x}\\ {\breve{v}}_{i}^{y}\\ \end{bmatrix},\ {\breve{A}}=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\ {\breve{B}}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&1\end{bmatrix}$,
${\breve{u}}_{i}=$ $\begin{bmatrix}\breve{a}_{i}^{x}\\ \breve{a}_{i}^{y}\end{bmatrix},\ {\breve{C}}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}$,
with ${\breve{p}}_{i}^{x}$ and ${\breve{p}}_{i}^{y}$ denoting the positions, ${\breve{v}}_{i}^{x}$ and ${\breve{v}}_{i}^{y}$ the velocities, and $\breve{a}_{i}^{x}$ and $\breve{a}_{i}^{y}$ the accelerations of the $i$th agent. Based on a connected graph $\mathcal{G}$ and a nonempty set $\mathcal{R}$, the agents collect relative/absolute input-output information, i.e., $u_{i}$ and $y_{i}$ defined in (12), and carry out the distributed omniscient observers presented in Theorem 2.
Leaders’ Actions: Leaders control their movements according to the estimated positions and velocities of themselves. In other words, a leader with index $j$ will use the estimate for ${\breve{x}}_{j}$ provided in $\hat{x}_{j}$ to design control input ${\breve{u}}_{j}$ for itself. If the leader intends to attract followers to position $(\bar{p}^{x}_{j},\bar{p}^{y}_{j})$, it will control itself to asymptotically move at a speed $v^{*}_{j}$ anticlockwise along the circumference of a circle centered at $(\bar{p}^{x}_{j},\bar{p}^{y}_{j})$ with a unit radius.
Followers’ Decisions: Each follower determines in real time a candidate leader for itself, based on the estimates of the positions and velocities of all agents produced by the distributed omniscient observers. From the perspective of a follower agent, each of other agents will be labeled as a candidate leader/follower, if the estimated speed (the Euclidean norm of the estimated velocity vector) of the agent is greater/less than ${v}^{*}_{t}$, a positive threshold chosen to be lower than the minimum circling speed of the leaders. Then, the follower will assign each candidate follower (including itself) a candidate leader that is as near as possible, such that the number of each candidate leader’s candidate followers is in proportion to the estimates of the candidate leaders’ speeds.
Followers’ Actions: Each follower heads toward a candidate target point indicated by its candidate leader. The point is one unit away from the estimated position of the candidate leader, along the direction indicated by a $\pi/2$ anticlockwise rotation of the estimated velocity vector of the candidate leader. A follower with index $k$ will use the estimate for ${\breve{x}}_{k}$ provided in $\hat{x}_{k}$ to design control input ${\breve{u}}_{k}$ for itself.
Figure 2: Snapshots of agent positions in Section V-B.
Figure 3: Agent positions vs. time in Section V-B.
V-C Simulation Results and Discussions
Figs. 1-3 demonstrate the simulation results of the two examples formulated in Section V-A and V-B, respectively. The agents that have access to their own positions are agents 2 and 7 in the first example, and agents 2, 8, and 11 in the second example. In addition, the three leaders’ circling speeds are designed to be $v^{*}_{3}=2$, $v^{*}_{6}=3$, and $v^{*}_{9}=4$. Other detailed settings are not listed here for the sake of brevity. The following are several interpretations and discussions of the simulation results.
  • In both examples, just a few agents have access to their own positions. Other agents only access relative position information. Moreover, the agents measure neither the absolute nor the relative velocity information.
  • Fig. 1 shows that all followers can move into the convex hull formed by the leaders. Since the leaders are allowed to move freely, the shape and position of the convex hull can be time-varying.
  • Fig. 2 shows that the leaders wiggle before circling. This is because they use the state estimate, not the true state, to control themselves.
  • Fig. 3 shows that two, three, and four followers are attracted to the three leaders respectively, in proportion to the circling speeds of the leaders. Different initial positions of the agents may lead to a different final result, but will preserve the ratio of the number of followers that each leader attracts.
  • In both examples, each follower can identify who are the leaders autonomously. This is achieved by estimating the augmented identity state (in the first example) and the velocity (in the second example) of agents.
  • Such leader-identification mechanisms can accommodate the change of leaders. Suppose in the second example that agents 3 and 6 keep circling at their speeds, while agent 9 stops circling after $t=20$s. Then, for agent 9 and its four followers, two of them will head toward position $(1,1)$, and the other three will go to $(5,5)$ autonomously.
VI Conclusion
In this article, distributed omniscient observers are proposed for heterogeneous and homogeneous linear MAS, respectively. The observer design for the latter is based on mostly relative, as well as a small amount of absolute, local input-output information of the agents. Knowledge of the global communication graph can obviate the relative input information required for the design, and vice versa. As a result, each agent can estimate the states of itself and other agents. The state estimation errors can converge to zero even though the inputs of other agents are persistently nonzero. An application in distributed Nash equilibrium seeking, and two bio-inspired simulation examples show that the proposed distributed omniscient observers can contribute to the emergence of collective intelligence in MAS.
VII Appendix
VII-A Proof of Lemma 2
First, consider the case where set $\mathcal{R}$ contains exactly one element $r$. There exists a permutation operation that rearranges the $r$th row and column in matrix ${\rm{col}}(\bar{\mathcal{L}}_{i})_{i=1}^{N}$ to the bottom and the rightmost respectively, i.e.,
$W^{\top}{\rm{col}}(\bar{\mathcal{L}}_{i})_{i=1}^{N}W=\begin{bmatrix}{\mathcal{L}}^{W}_{1,1}&*\\ 0&1\end{bmatrix}$,
where $W$ is a permutation matrix, and ${\mathcal{L}}^{W}_{1,1}$ is a nonsingular matrix according to Lemma 5 in [3]. It follows that ${\rm{col}}(\bar{\mathcal{L}}_{i})_{i=1}^{N}$ is also nonsingular. For the case where set $\mathcal{R}$ contains more than one element, similar analysis can be done.
VII-B Proof of Lemma 3
According to (13) and (14c),
$T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)\!=\!\begin{cases}{\left\|{{\bar{\cal L}}_{i}}\right\|^{-1}}\!\begin{bmatrix}l_{i1}\breve{B}&l_{i2}\breve{B}&\cdots&l_{iN}\breve{B}\end{bmatrix},\ i\in\mathcal{N}\setminus\mathcal{R}\\[5.0pt] {\left\|{{\bar{\cal L}}_{i}}\right\|^{-1}}\!\begin{bmatrix}0_{\breve{n}\times(i-1)\breve{m}}&\breve{B}&0_{\breve{n}\times(N-i)\breve{m}}\end{bmatrix},\ i\in\mathcal{R}.\end{cases}$
Therefore, ${\rm{Im}}\breve{B}={\rm{Im}}\left[{T_{id}^{\top}\left({{I_{N}}\otimes\breve{B}}\right)}\right]$ and Lemma 3 follows.
VII-C Proof of Theorem 3
It follows from [29] that $\left\|{{{\hat{\breve{x}}}_{i}}(t)-{\breve{x}}_{i}(t)}\right\|$, i.e., $\left\|T_{id}^{\top}\left(\hat{x}_{i}(t)-x(t)\right)\right\|$, exponentially converges to zero, $\forall i\in\mathcal{N}$. Then it can be proved as in [4] that adaptive gains remain bounded and $\mathop{\lim}\nolimits_{t\to\infty}\left\|T_{iu}^{\top}\left(\hat{x}_{i}(t)-x(t)\right)\right\|=0,\ \forall i\in\mathcal{N}$.
VII-D Proof of Theorem 4
Consider the following Lyapunov candidate function
${V}=\frac{1}{2}\left(x-x^{*}\right)^{\top}\left(x-x^{*}\right)+V_{o},$ (24)
where $V_{o}$ is a Lyapunov function constructed in [4], that is
$V_{o}$ $=\frac{1}{2}\varepsilon_{u}^{\top}{\left[{T_{u}^{\top}({\cal L}\otimes I){T_{u}}}\right]^{-1}}{\varepsilon_{u}}+\sum\nolimits_{i=1}^{N}{\frac{1}{{2{\phi_{i}}}}{{({\gamma_{i}}-{\gamma^{*}})}^{2}}}$
$\ \ \ \ +\sum\nolimits_{i=1}^{N}{\frac{1}{{2{\phi_{is}}}}{{({\gamma_{is}}-\gamma_{s}^{*})}^{2}}}.$ (25)
In (25), $\varepsilon_{u}={\rm col}\left(\varepsilon_{iu}\right)_{i=1}^{N}$, $\varepsilon_{iu}=T_{iu}^{\top}\sum\nolimits_{j=1}^{N}{a_{ij}(\hat{x}_{i}-\hat{x}_{j})}$, $T_{u}={\rm col}\left(T_{iu}\right)_{i=1}^{N}$, and $\gamma^{*}$ and $\gamma_{s}^{*}$ are two positive constants to be determined. Differentiating (24) along the trajectory of (21) yields that
$\dot{V}=$ $-\left(x-x^{*}\right)^{\top}{\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)+\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})-\nabla_{\breve{x}_{i}}J_{i}(x)\right]_{i=1}^{N}+\dot{V}_{o}$
$\leq$ $-\mu\left\|x-x^{*}\right\|^{2}\!+\left(x-x^{*}\right)^{\top}\!\!{\rm{col}}\!\left[\nabla_{\breve{x}_{i}}J_{i}(x)-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})\right]_{i=1}^{N}\!\!+\dot{V}_{o}$
$\leq$ $-\frac{\mu}{2}\left\|x-x^{*}\right\|^{2}+\frac{1}{2\mu}\left\|{\rm{col}}\left[\nabla_{\breve{x}_{i}}J_{i}(x)-\nabla_{\breve{x}_{i}}J_{i}(\hat{x}_{i})\right]_{i=1}^{N}\right\|^{2}+\dot{V}_{o}$
$\leq$ $-\frac{\mu}{2}\left\|x-x^{*}\right\|^{2}+V_{o}^{\prime},$
where $V_{o}^{\prime}=\frac{\chi}{2\mu}\sum\nolimits_{i=1}^{N}\left\|x-\hat{x}_{i}\right\|^{2}+\frac{\chi_{s}}{2\mu}\sum\nolimits_{i=1}^{N}\left\|x-\hat{x}_{i}\right\|+\dot{V}_{o}$. It can be verified, by using the same analysis method as in [4], that there exist four positive constants $\gamma^{*}$, $\gamma_{s}^{*}$, $\lambda^{*}$, and $\lambda^{*}_{s}$, such that $V_{o}^{\prime}\leq\lambda^{*}\left\|{{\varepsilon_{d}}}\right\|^{2}+\lambda^{*}_{s}\left\|{{\varepsilon_{d}}}\right\|$, where $\varepsilon_{d}={\rm col}\left(\varepsilon_{id}\right)_{i=1}^{N}$ and ${\varepsilon_{id}}=T_{id}^{\top}\left({\hat{x}}_{i}-x\right)$. Therefore,
$\dot{V}-\lambda^{*}\left\|{{\varepsilon_{d}}}\right\|^{2}-\lambda^{*}_{s}\left\|{{\varepsilon_{d}}}\right\|\leq-\frac{\mu}{2}\left\|x-x^{*}\right\|^{2}.$
Similar to the proof in [4], it can be proved that $V$ is bounded, and therefore, $x$, $\hat{x}_{i}$, $\gamma_{i}$, and $\gamma_{is}$ are all bounded. According to (21) and (22), $\dot{x}$ is also bounded, which guarantees that $x$ is uniformly continuous. Then it follows from the Barbalat’s Lemma [15] that $\mathop{\lim}\nolimits_{t\to\infty}\left\|x(t)-x^{*}\right\|=0$.